# LLM-in-Sandbox Elicits General Agentic Intelligence

### v2

@@ 2.4.2 Quantitative Analysis
Table 4: Sandbox capability usage rate comparison: strong models (average of all models except Qwen3-4B-Instruct) vs. weak model (Qwen3-4B-Instruct). Usage rate = capability invocations / total turns.
3 LLM-in-Sandbox Reinforcement Learning Enhances Generalization
The preceding experiments demonstrate that the sandbox environment holds significant potential for enhancing general intelligence: strong agentic models consistently benefit across diverse domains. This raises a natural question: can we directly train LLMs within the sandbox to further unlock their potential? Motivated by this, we propose LLM-in-Sandbox Reinforcement Learning (LLM-in-Sandbox-RL), which trains models on general context-based tasks within the sandbox, enabling them to effectively explore the sandbox environment without requiring specialized agentic data.
3.1 Method
To train LLMs to effectively utilize code sandboxes for general tasks, the ideal approach should satisfy two criteria: (1) training within a sandbox to learn to explore the environment, and (2) using general-domain data to ensure broad transferability. As shown in Table 5, existing approaches achieve some but not all of these goals. For example, vanilla RL training for LLMs on text-only tasks (hereafter LLM-RL; Lambert et al., 2024), though capable of leveraging general data, does not involve sandbox interaction; SWE-RL (Wei et al., 2025; Luo et al., 2025), which trains models on software engineering tasks within sandboxes, enables sandbox interaction but relies on domain-specific data.
We propose training LLMs within sandbox environments configured with general-purpose data. Specifically, we adopt context-based tasks: each task consists of background materials (e.g., documents) and a task objective that must be completed based on these materials. Since completing the objective depends on the provided materials, models must actively explore the sandbox to find relevant information, naturally learning to leverage its capabilities. Meanwhile, our method uses only a general-purpose sandbox without task-specific configurations, making it easy to scale. Moreover, general-purpose tasks are much simpler to curate than software engineering tasks, enabling easy data scaling (Cheng et al., 2024; Jain et al., 2025). Overall, LLM-in-Sandbox-RL combines the benefits of sandbox-based training with general-domain data and easy scalability.
LLM-RL SWE-RL LLM-in-Sandbox-RL
Sandbox Utilization
General Domain
Data Scalability
Environment Scalability N/A
Table 5: Comparison of different RL training paradigms. LLM-RL refers to RL training for LLMs on text-only tasks; SWE-RL refers to RL training on software engineering tasks within sandboxes.
Data Source.
Sandbox Configuration.
As illustrated in Figure 3, we store task contexts as files within the sandbox environment. To enrich the context and increase the task difficulty, we employ several strategies:
Figure 3: Sandbox Configuration for LLM-in-Sandbox-RL: task contexts are stored as files within the sandbox environment. (a) Multi-document or long contexts are split into separate files. (b) Single-file contexts are supplemented with distractors.
Task Setup.
RL Training.
Recent advances in RL for LLMs typically adopt outcome-based rewards: given a prompt, the model generates a trajectory, and the entire trajectory is rewarded based on the correctness of the final output (Guo et al., 2025). Our RL baseline follows this paradigm to train LLMs on context-based tasks without sandbox interaction, which we refer to as LLM-RL. LLM-in-Sandbox-RL adopts the same framework, with the only difference being that trajectory generation uses LLM-in-Sandbox mode instead of vanilla LLM mode.
3.2 Experiments
Setup
Math Physics Chem. Biomed. Long-Cont. Instruct. SWE
LLM LLM LLM LLM LLM LLM LLM LLM LLM LLM LLM LLM LLM
Qwen3-4B-Instruct-2507
Base LLM 46.0 32.5 22.9 25.6 56.4 49.3 10.2 9.4 8.6 10.5 33.7 29.0 11.2
LLM-RL 48.5 33.1 21.3 26.0 52.4 50.4 9.4 12.8 9.1 14.1 35.3 34.0 12.8
$\Delta$ +2.5 +0.6 -1.6 +0.4 -4.0 +1.1 -0.8 +3.4 +0.5 +3.6 +1.7 +5.0 +1.6
LLM-RL 51.0 53.3 23.0 30.3 57.8 63.3 10.8 13.2 10.6 17.6 33.7 38.7 12.8
$\Delta$ +5.0 +20.8 +0.1 +4.7 +1.3 +14.0 +0.6 +3.8 +2.0 +7.1 0.0 +9.7 +1.6
Qwen3-Coder-30B-A3B
Base LLM 26.0 41.5 23.9 28.9 48.7 54.0 14.4 18.2 24.6 23.9 34.7 40.3 45.0
LLM-RL 28.1 40.9 23.9 29.5 50.5 57.1 16.8 15.6 26.5 21.9 34.0 39.3 47.6
$\Delta$ +2.1 -0.6 0.0 +0.6 +1.8 +3.1 +2.4 -2.6 +1.9 -2.0 -0.7 -1.0 +2.6
LLM-RL 29.8 42.7 25.4 29.3 49.1 56.7 18.6 17.6 24.6 36.1 31.7 38.0 48.0
$\Delta$ +3.8 +1.2 +1.5 +0.4 +0.4 +2.7 +4.2 -0.6 0.0 +12.2 -3.0 -2.3 +3.0
Main Results.
Data Source and Context Placement
To understand the impact of training data, we compare four variants of LLM-in-Sandbox-RL on Qwen3-4B-Instruct-2507: (1) Math: mathematical data from DAPO (Yu et al., 2025); (2) SWE: software engineering data from R2E-Gym (Jain et al., 2025); (3) Gen. in Prompt: our general context-based data with context placed in the prompt; and (4) Gen. in Sandbox: the same data but with context placed in the sandbox. As shown in Table 7, all variants achieve some degree of cross-domain generalization, demonstrating the broad applicability of our training paradigm. Among them, Gen. Sandbox achieves the best overall performance. Notably, the comparison between Gen. Prompt and Gen. Sandbox highlights the importance of sandbox interaction: placing context in the sandbox forces the model to actively explore the environment, yielding stronger generalization than directly providing context in the prompt.
Math Physics Chem. Biomed. Long-Cont. Instruct. SWE
LLM LLM LLM LLM LLM LLM LLM LLM LLM LLM LLM LLM LLM
Base LLM 46.0 32.5 22.9 25.6 56.4 49.3 10.2 9.4 8.6 10.5 33.7 29.0 11.2
Math 53.5 51.0 23.4 31.4 58.2 60.0 9.6 13.6 9.5 11.3 33.0 37.0 15.2
SWE 47.5 36.5 21.3 21.2 57.8 47.8 9.4 9.2 11.6 9.9 33.0 33.0 17.4
3.3 Analysis on Generalization
To understand why LLM-in-Sandbox-RL training leads to broad generalization, we first examine how models change their sandbox capability usage after training, and then investigate how sandbox-mode training transfers to LLM mode. We adopt the same capability classification and quantification method as in Section 2.4.
Generalization across Domains.
Table 8 shows that after training, models exhibit increased sandbox capability usage across all three dimensions (external resources, file management, computation). As analyzed in Section 2, these capabilities benefit diverse task domains, explaining why the learned exploration skills transfer broadly.
Model External File Computation Avg. Turns
Qwen3-4B-Instruct-2507 Base LLM 0.8% 2.9% 2.9% 23.7
LLM-RL 4.1% 7.3% 7.2% 7.0
Qwen3-Coder-30B-A3B Base LLM 5.7% 24.1% 11.1% 9.5
LLM-RL 5.7% 24.4% 11.9% 10.0
Generalization across Model Capabilities.
As shown in Table 8, weaker models (Qwen3-4B) show larger improvements: capability usage rate increases substantially while average turns decrease dramatically from 23.7 to 7.0. Recall from Section 2.4.2 that the base Qwen3-4B model “wanders” in the sandbox with many ineffective turns; after LLM-in-Sandbox-RL training, the model learns to accomplish tasks with fewer but more purposeful interactions. Stronger models (Qwen3-Coder) already have high capability usage, so improvements are more modest but still consistent.
Generalization across Inference Modes.
We analyze reasoning patterns in outputs of vanilla LLM mode before and after LLM-in-Sandbox-RL training. Specifically, we measure two categories of behaviors through pattern matching: (1) Structural Organization: Markdown formatting elements (headers, separators, bullet points, math blocks) that indicate explicit step-by-step reasoning; and (2) Verification Behaviors: phrases indicating self-checking (e.g., “let’s verify”, “check that”, confirmation markers). As shown in Table 9, both models exhibit increased structural organization and verification behaviors after training. These reasoning patterns, learned through multi-turn sandbox interaction where each action receives explicit feedback, transfer to LLM mode even without sandbox access.
Model Verification Structure
Base LLM LLM-RL Base LLM LLM-RL
Qwen3-4B-Instruct-2507 20.22 36.91 19.13 20.64
Qwen3-Coder-30B-A3B 0.77 0.88 10.30 16.12
4 LLM-in-Sandbox Enables Efficient Deployment
@@ 4.2 Sandbox Infrastructure
5 LLM-in-Sandbox Goes Beyond Text Generation
@@
Previous sections evaluate LLM-in-Sandbox on tasks where both vanilla LLMs and LLM-in-Sandbox can produce outputs for end-to-end comparison. However, LLM-in-Sandbox also enables capabilities that are fundamentally impossible for standalone LLMs. By granting LLMs access to a virtual computer, LLM-in-Sandbox transcends the text-in-text-out paradigm and unlocks new possibilities:
@@ Discussion.
Nevertheless, these cases reveal a promising direction: as LLMs become more capable and sandbox environments more sophisticated, LLM-in-Sandbox could evolve into a truly general-purpose digital creation system. We believe this paradigm, LLMs interacting with computational environments rather than generating text in isolation, represents a compelling path toward general intelligence.
@@ 6 Conclusion and Future Work
LLM-in-Sandbox as Default Inference Infrastructure
We introduce LLM-in-Sandbox, a paradigm that grants LLMs access to a virtual computer and show that strong LLMs exhibit emergent capabilities to leverage this environment for general tasks. Looking forward, we envision LLM-in-Sandbox becoming the default paradigm for serving LLMs: analytical tasks gain verifiable computation, long-context tasks benefit from file-based management, and creative tasks yield actual outputs (e.g., images, videos and applications) rather than text descriptions. We anticipate sandbox environments will become standard infrastructure, transforming LLMs from text generators into general-purpose digital workers.
LLM-in-Sandbox as an Agentic Capability Benchmark
Beyond serving as an inference paradigm, LLM-in-Sandbox naturally provides a standardized testbed for evaluating agentic capabilities. Unlike existing benchmarks that focus on specific downstream tasks, LLM-in-Sandbox measures fundamental skills like exploration, tool use, and self-verification through a unified framework. The metric $\Delta=\text{LLM-in-Sandbox}-\text{LLM}$ offers a meaningful indicator: it quantifies how effectively a model can leverage computational environments, revealing agentic potential that raw LLM performance alone cannot capture.
Sandbox-Native Model Training
We propose LLM-in-Sandbox-RL, a lightweight RL method that trains sandbox interaction as a transferable skill using only general, non-agentic data. Looking ahead, we advocate for sandbox-native models where sandbox interaction becomes a first-class training objective, not only through large-scale RL with real environmental feedback, but also by incorporating sandbox-style reasoning into the pretraining stage itself.
@@ Appendix A Sandbox Implementation
execute_bash Description: Execute a bash command in the terminal within a persistent shell session. One command at a time: You can only execute one bash command at a time. If you need to run multiple commands sequentially, use && or ; to chain them together. Persistent session: Commands execute in a persistent shell session where environment variables, virtual environments, and working directory persist between commands. Soft timeout: Commands have a soft timeout of 10 seconds, once that’s reached, you have the option to continue or interrupt the command. Output truncation: If the output exceeds a maximum length, it will be truncated before being returned. Parameters: command (string, required): The bash command to execute. For example: python my_script.py.
str_replace_editor Description: Custom editing tool for viewing, creating and editing files. State is persistent across command calls and discussions with the user. If path is a file, view displays the result of applying cat -n. If path is a directory, view lists non-hidden files and directories up to 2 levels deep. The create command cannot be used if the specified path already exists as a file. For the str_replace command, the old_str parameter should match EXACTLY one or more consecutive lines from the original file. Parameters: command (string, required): The command to run. Allowed options are: view, create, str_replace, insert. path (string, required): Absolute path to file or directory. file_text (string, optional): Required for create command. old_str (string, optional): Required for str_replace command. new_str (string, optional): The replacement string for str_replace, or the string to insert for insert. insert_line (integer, optional): Required for insert command. view_range (array, optional): Line range for view command, e.g., [11, 12].
submit Description: Finish the interaction when the task is complete OR if the assistant cannot proceed further with the task. Parameters: No parameters are required for this function.
@@ Appendix C Evaluation Details
We evaluate on six diverse non-code domains, summarized in Table 14. The system prompt largely follows the one in Appendix F, with minor domain-specific adjustments and final answer formatting instructions. Please refer to our released code for the exact prompts.
@@
Physics UGPhysics (Xu et al., ) 650 LLM Judge
@@
Biomedicine MedXpertQA (Zuo et al., ) 500 Exact Match
Long-Context AA-LCR (Team, 2025) 100 $\times$ 4 LLM Equality Checker
@@
Appendix D Sandbox Capability Classification

### v3

@@ 2.4.2 Quantitative Analysis
Table 4: Computer capability usage rate comparison: strong models (average of all models except Qwen3-4B-Instruct) vs. weak model (Qwen3-4B-Instruct). Usage rate = capability invocations / total turns.
3 Computer Environments Enable Efficient Deployment
@@ 3.2 Sandbox Infrastructure
4 Reinforcement Learning within Computers Enhances Generalization
The experiments in Section 2 demonstrate that the computer environment holds significant potential for enhancing general intelligence: strong agentic models consistently benefit across diverse domains. This raises a natural question: can we directly train LLMs within the computer environment to further unlock their potential? Motivated by this, we propose LLM-in-Sandbox Reinforcement Learning (LLM-in-Sandbox-RL), which trains models on general context-based tasks within the computer environment, enabling them to effectively explore the environment without requiring specialized agentic data.
4.1 Method
To train LLMs to effectively utilize computer environments for general tasks, the ideal approach should satisfy two criteria: (1) training within the computer environment to learn to explore the environment, and (2) using general-domain data to ensure broad transferability. As shown in Table 8, existing approaches achieve some but not all of these goals. For example, vanilla RL training for LLMs on text-only tasks (hereafter LLM-RL; Lambert et al., 2024), though capable of leveraging general data, does not involve computer environment interaction; SWE-RL (Wei et al., 2025b; Luo et al., 2025), which trains models on software engineering tasks within computer environments, enables environment interaction but relies on domain-specific data.
We propose training LLMs within computer environments configured with general-purpose data. Specifically, we adopt context-based tasks: each task consists of background materials (e.g., documents) and a task objective that must be completed based on these materials. Since completing the objective depends on the provided materials, models must actively explore the computer environment to find relevant information, naturally learning to leverage its capabilities. Meanwhile, our method uses only a general-purpose computer environment without task-specific configurations, making it easy to scale. Moreover, general-purpose tasks are much simpler to curate than software engineering tasks, enabling easy data scaling (Cheng et al., 2024; Jain et al., 2025). Overall, LLM-in-Sandbox-RL combines the benefits of computer environment-based training with general-domain data and easy scalability.
LLM-RL SWE-RL LLM-in-Sandbox-RL
Computer Utilization
General Domain
Data Scalability
Environment Scalability N/A
Table 8: Comparison of different RL training paradigms. LLM-RL refers to RL training for LLMs on text-only tasks; SWE-RL refers to RL training on software engineering tasks within computer environments.
Data Source.
Computer Environment Configuration.
As illustrated in Figure 3, we store task contexts as files within the computer environment. To enrich the context and increase the task difficulty, we employ several strategies:
Figure 3: Environment Configuration for LLM-in-Sandbox-RL: task contexts are stored as files within the sandbox. (a) Multi-document or long contexts are split into separate files. (b) Single-file contexts are supplemented with distractors.
Task Setup.
RL Training.
Recent advances in RL for LLMs typically adopt outcome-based rewards: given a prompt, the model generates a trajectory, and the entire trajectory is rewarded based on the correctness of the final output (Guo et al., 2025). Our RL baseline follows this paradigm to train LLMs on context-based tasks without environmental interaction, which we refer to as LLM-RL. LLM-in-Sandbox-RL adopts the same framework, with the only difference being that trajectory generation uses LLM-in-Sandbox mode instead of vanilla LLM mode.
4.2 Experiments
Setup
Math Physics Chem. Biomed. Long-Cont. Instruct. SWE
LLM LLM LLM LLM LLM LLM LLM LLM LLM LLM LLM LLM LLM
Qwen3-4B-Instruct-2507
Base LLM 46.0 32.5 22.9 25.6 56.4 49.3 10.2 9.4 8.6 10.5 33.7 29.0 11.2
LLM-RL 48.5 33.1 21.3 26.0 52.4 50.4 9.4 12.8 9.1 14.1 35.3 34.0 12.8
$\Delta$ +2.5 +0.6 -1.6 +0.4 -4.0 +1.1 -0.8 +3.4 +0.5 +3.6 +1.7 +5.0 +1.6
LLM-RL 51.0 53.3 23.0 30.3 57.8 63.3 10.8 13.2 10.6 17.6 33.7 38.7 12.8
$\Delta$ +5.0 +20.8 +0.1 +4.7 +1.3 +14.0 +0.6 +3.8 +2.0 +7.1 0.0 +9.7 +1.6
Qwen3-Coder-30B-A3B
Base LLM 26.0 41.5 23.9 28.9 48.7 54.0 14.4 18.2 24.6 23.9 34.7 40.3 45.0
LLM-RL 28.1 40.9 23.9 29.5 50.5 57.1 16.8 15.6 26.5 21.9 34.0 39.3 47.6
$\Delta$ +2.1 -0.6 0.0 +0.6 +1.8 +3.1 +2.4 -2.6 +1.9 -2.0 -0.7 -1.0 +2.6
LLM-RL 29.8 42.7 25.4 29.3 49.1 56.7 18.6 17.6 24.6 36.1 31.7 38.0 48.0
$\Delta$ +3.8 +1.2 +1.5 +0.4 +0.4 +2.7 +4.2 -0.6 0.0 +12.2 -3.0 -2.3 +3.0
Main Results.
Data Source and Context Placement
To understand the impact of training data, we compare four variants of LLM-in-Sandbox-RL on Qwen3-4B-Instruct-2507: (1) Math: mathematical data from DAPO (Yu et al., 2025); (2) SWE: software engineering data from R2E-Gym (Jain et al., 2025); (3) Gen. in Prompt: our general context-based data with context placed in the prompt; and (4) Gen. in Sandbox: the same data but with context placed in the sandbox. As shown in Table 10, all variants achieve some degree of cross-domain generalization, demonstrating the broad applicability of our training paradigm. Among them, Gen. Sandbox achieves the best overall performance. Notably, the comparison between Gen. Prompt and Gen. Sandbox highlights the importance of environmental interaction: placing context in the sandbox forces the model to actively explore the environment, yielding stronger generalization than directly providing context in the prompt.
Math Physics Chem. Biomed. Long-Cont. Instruct. SWE
LLM LLM LLM LLM LLM LLM LLM LLM LLM LLM LLM LLM LLM
Base LLM 46.0 32.5 22.9 25.6 56.4 49.3 10.2 9.4 8.6 10.5 33.7 29.0 11.2
Math 53.5 51.0 23.4 31.4 58.2 60.0 9.6 13.6 9.5 11.3 33.0 37.0 15.2
SWE 47.5 36.5 21.3 21.2 57.8 47.8 9.4 9.2 11.6 9.9 33.0 33.0 17.4
4.3 Analysis on Generalization
To understand why LLM-in-Sandbox-RL training leads to broad generalization, we first examine how models change their computer capability usage after training, and then investigate how sandbox-mode training transfers to LLM mode. We adopt the same capability classification and quantification method as in Section 2.4.
Generalization across Domains.
Table 11 shows that after training, models exhibit increased computer capability usage across all three dimensions (external resources, file management, computation). As analyzed in Section 2, these capabilities benefit diverse task domains, explaining why the learned exploration skills transfer broadly.
Model External File Computation Avg. Turns
Qwen3-4B-Instruct-2507 Base LLM 0.8% 2.9% 2.9% 23.7
LLM-RL 4.1% 7.3% 7.2% 7.0
Qwen3-Coder-30B-A3B Base LLM 5.7% 24.1% 11.1% 9.5
LLM-RL 5.7% 24.4% 11.9% 10.0
Generalization across Model Capabilities.
As shown in Table 11, weaker models (Qwen3-4B) show larger improvements: capability usage rate increases substantially while average turns decrease dramatically from 23.7 to 7.0. Recall from Section 2.4.2 that the base Qwen3-4B model “wanders” in the computer environment with many ineffective turns; after LLM-in-Sandbox-RL training, the model learns to accomplish tasks with fewer but more purposeful interactions. Stronger models (Qwen3-Coder) already have high capability usage, so improvements are more modest but still consistent.
Generalization across Inference Modes.
We analyze reasoning patterns in outputs of vanilla LLM mode before and after LLM-in-Sandbox-RL training. Specifically, we measure two categories of behaviors through pattern matching: (1) Structural Organization: Markdown formatting elements (headers, separators, bullet points, math blocks) that indicate explicit step-by-step reasoning; and (2) Verification Behaviors: phrases indicating self-checking (e.g., “let’s verify”, “check that”, confirmation markers). As shown in Table 12, both models exhibit increased structural organization and verification behaviors after training. These reasoning patterns, learned through multi-turn environmental interaction where each action receives explicit feedback, transfer to LLM mode even without environment access.
Model Verification Structure
Base LLM LLM-RL Base LLM LLM-RL
Qwen3-4B-Instruct-2507 20.22 36.91 19.13 20.64
Qwen3-Coder-30B-A3B 0.77 0.88 10.30 16.12
5 LLMs within Computer Environments Go Beyond Text Generation
@@
Previous sections evaluate LLM-in-Sandbox on tasks where both vanilla LLMs and LLM-in-Sandbox can produce outputs for end-to-end comparison. However, LLM-in-Sandbox also enables capabilities that are fundamentally impossible for standalone LLMs. By granting LLMs access to the computer environment, LLM-in-Sandbox transcends the text-in-text-out paradigm and unlocks new possibilities:
@@ Discussion.
Nevertheless, these cases reveal a promising direction: as LLMs become more capable and computer environments more sophisticated, LLM-in-Sandbox could evolve into a truly general-purpose digital creation system. We believe this paradigm, LLMs interacting with computational environments rather than generating text in isolation, represents a compelling path toward general intelligence.
@@ 6 Conclusion and Future Work
Computer Environments as Default Inference Infrastructure
We introduce LLM-in-Sandbox, a paradigm that grants LLMs access to a virtual computer and show that strong LLMs exhibit emergent capabilities to leverage this environment for general tasks. Looking forward, we envision LLM-in-Sandbox becoming the default paradigm for serving LLMs: analytical tasks gain verifiable computation, long-context tasks benefit from file-based management, and creative tasks yield actual outputs (e.g., images, videos and applications) rather than text descriptions. We anticipate computer environments will become standard infrastructure, transforming LLMs from text generators into general-purpose digital workers.
Computer Environments as an Agentic Capability Benchmark
Beyond serving as an inference paradigm, LLM-in-Sandbox naturally provides a standardized testbed for evaluating agentic capabilities. Unlike existing benchmarks that focus on specific downstream tasks (Jimenez et al., 2023; Yao et al., 2025; Wei et al., 2025a), LLM-in-Sandbox measures fundamental skills like exploration, tool use, and self-verification through a unified framework. The metric $\Delta=\text{LLM-in-Sandbox}-\text{LLM}$ offers a meaningful indicator: it quantifies how effectively a model can leverage computational environments, revealing agentic potential that raw LLM performance alone cannot capture.
Computer-Native Model Training
We propose LLM-in-Sandbox-RL, a lightweight RL method that trains computer interaction as a transferable skill using only general, non-agentic data. Looking ahead, we advocate for computer-native models where computer interaction becomes a first-class training objective, not only through large-scale RL with real environmental feedback, but also by incorporating computer-style reasoning into the pretraining stage itself.
@@ Appendix A Sandbox Implementation
bash Description: Execute a bash command in the terminal within a persistent shell session. One command at a time: You can only execute one bash command at a time. If you need to run multiple commands sequentially, use && or ; to chain them together. Persistent session: Commands execute in a persistent shell session where environment variables, virtual environments, and working directory persist between commands. Soft timeout: Commands have a soft timeout of 10 seconds, once that’s reached, you have the option to continue or interrupt the command. Output truncation: If the output exceeds a maximum length, it will be truncated before being returned. Parameters: command (string, required): The bash command to execute. For example: python my_script.py.
file_editor Description: Custom editing tool for viewing, creating and editing files. State is persistent across command calls and discussions with the user. If path is a file, view displays the result of applying cat -n. If path is a directory, view lists non-hidden files and directories up to 2 levels deep. The create command cannot be used if the specified path already exists as a file. For the str_replace command, the old_str parameter should match EXACTLY one or more consecutive lines from the original file. Parameters: command (string, required): The command to run. Allowed options are: view, create, str_replace, insert. path (string, required): Absolute path to file or directory. file_text (string, optional): Required for create command. old_str (string, optional): Required for str_replace command. new_str (string, optional): The replacement string for str_replace, or the string to insert for insert. insert_line (integer, optional): Required for insert command. view_range (array, optional): Line range for view command, e.g., [11, 12].
finish Description: Finish the interaction when the task is complete OR if the assistant cannot proceed further with the task. Parameters: No parameters are required for this function.
@@ Appendix C Evaluation Details
We evaluate on six non-code domains and one code domain, summarized in Table 14. The system prompt largely follows the one in Appendix F, with minor domain-specific adjustments and final answer formatting instructions. Please refer to our released code for the exact prompts.
@@
Physics UGPhysics (Xu et al., 2025) 650 LLM Judge
@@
Biomedicine MedXpertQA (Zuo et al., 2025) 500 Exact Match
Long-Context AA-LCR (Artificial Analysis Team, 2025) 100 $\times$ 4 LLM Equality Checker
@@
Appendix D Computer Capability Classification